[figure: 9.7 cm]

MAGNETOVOLUME EFFECT IN PARAMAGNETIC YbAgCu_{4}

COMPOUND

I.V. Svechkarev^{a}, A.S. Panfilov^{a}, S.N. Dolja^{a}, H. Nakamura^{b}, M. Shiga^{b}

aB. Verkin Institute for Low Temperature Physics, Kharkov, 310164, Ukraine bDepartment of Materials Science and Engineering, Kyoto University Kyoto 606-01, Japan

Magnetovolume effect dlnχ/dlnV in the YbAgCu_{4} compound with weak intermediate valence has been studied in the temperature range 4.2–300 K. The low temperature paramagnetostriction and direct pressure dependence of the magnetic susceptibility χ at high temperatures were used. Grüneisen parameter for the characteristic Kondo temperature G_{K}=–dlnT_{K}/dlnV= –28 follows from the data obtained in Coqblin-Schriffer limit and changes slightly to –22 as temperature increases. This value is consistent with resistivity vs pressure data and is close to G_{K} in the YbInCu_{4} compound. More sophisticated treatment of the magnetovolume effect in resonant level model for susceptibility suggests that the shift of the resonant level rather than its width change under pressure is the dominant contribution to this effect.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : I.V. Svechkarev

Address for correspondence : B. Verkin Institute for Low Temperature Physics, 47, Lenin’s pr., Kharkov, 310164, Ukraine

Email address : isvechkarev@ilt.kharkov.ua